OpenStudy (nataliareina):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3).

OpenStudy (anonymous):

change your equation first to slope intercept form... what is the slope...?

OpenStudy (anonymous):

y = −two thirdsx − 1

OpenStudy (anonymous):

note: the slope of a line perpendicular to another equation is \[-(\frac{ 1 }{ m })\] negative of the reciprocal of m...

OpenStudy (anonymous):

As jesi1201 already told u how to find the gradient! U can use this formula to find the equation: y-y1= m(x-x1) where y1= 3 and x1= -6 btw 'm' is the gradient of the line :)